An electric immersion heater has a power rating of 1500 W. If the heater is placed in a liter of water at 20
1
Expert's answer
2021-08-16T08:35:42-0400

Here's what I believe the complete condition of the problem is:

An electric immersion heater has a power rating of 1500 W. If the heater is placed in a liter of water at 20°C, how many minutes will it take to bring the water to a boil? (Assume that there is no heat loss except to the water itself.)


So, the energy required to increase the temperature is


"Q=cm\Delta \tau."


The electrical energy required for this is equivalent to power times time:


"E=Pt,\\
Q=E,\\
cm\Delta\tau=Pt,\\\space\\
t=\frac{cm\Delta\tau}{P}=\frac{4200·1·(100-20)}{1500}=224\text{ s},"

or 3 min 44 s.